# ReaCritic: Large Reasoning Transformer-based DRL Critic-model Scaling For Heterogeneous Networks

### v1

@@ V Numerical Results
V-A Experimental Setup
V-A1 Wireless Benchmark Environment
We construct a high-dimensional wireless environment simulating a HetNet with $M\in\left[10,50\right]$ users. Each user is randomly placed within a normalized 2D area and exhibits independent mobility, modeled via Gaussian perturbations per time step. Channel conditions follow a Rician fading model combined with log-distance path loss, with transmission distances sampled from $\left[5,10\right]$ meters. Channel gains are updated at each episode to reflect temporal variation.
Each user is characterized by heterogeneous attributes, i.e., task demand, latency requirement, computation capability, channel quality, physical location, and a one-hot encoded type indicator. The maximum system capacities are set to $20$ Hz bandwidth, $10$ W transmission power, and $100$ computation units. At each decision step, the agent selects a five-dimensional action vector per user, controlling uplink/downlink power, bandwidth, and compute allocation. Actions are normalized and clipped to satisfy global resource constraints. The environment computes uplink and downlink rates, total task latency, and a multi-objective reward incorporating throughput, energy efficiency, resource usage, and latency satisfaction. This setting captures the key challenges of resource allocation in dynamic, multi-user HetNets.
V-A2 DRL Benchmark Environments
To evaluate the general applicability of ReaCritic, we conduct experiments across two categories of benchmark environments. First, we compare against standard DRL algorithms, i.e., SAC, DDPG, A3C, TD3, and PPO, within our custom wireless resource allocation environment. These serve as strong baselines for measuring learning stability, sample efficiency, and final policy performance. Second, to assess task generalizability beyond wireless systems, we integrate ReaCritic into these DRL algorithms and test them on three representative OpenAI Gym [46] control tasks with varying complexity:
@@ V-A2 DRL Benchmark Environments
These environments provide diverse and progressively challenging domains to validate ReaCritic’s ability to enhance value estimation and policy learning across different settings.
V-B Experiments Performance Analysis
V-B1 Performance Evaluation in HetNet with Varying User Density
@@
Figure 5: Comparison of training performance with block number to be $4$, and user number to be 10 across ReaCritic-based SAC method with varying numbers of HRea steps and VRea steps, and standard SAC. The final reward ReaCritic-based SAC with different settings of HRea step and VRea step for 10 users is also included.
@@
Figure 6: Comparison of training performance with block number to be $4$, and user number to be 20 across ReaCritic-based SAC method, and standard SAC. The final reward ReaCritic-based SAC with different settings of HRea step and VRea step for 20 users is also included.
@@
Figure 7: Comparison of training performance with block number to be $4$, and user number to be 50 across ReaCritic-based SAC method, and standard SAC. The final reward ReaCritic-based SAC with different settings of HRea step and VRea step for 50 users is also included.
Fig. 5 compares the training performance of ReaCritic-based SAC under different HRea and VRea configurations with a small HetNet of $10$ users. Compared to the standard SAC baseline, ReaCritic consistently achieves higher and more stable cumulative rewards. When both HRea and VRea steps are small, e.g., $(4,3)$ or $(8,1)$, the model already outperforms standard SAC. Moreover, the critic output remains stable and well-aligned with the reward trajectory, indicating a coherent value function approximation. In contrast, standard SAC shows unstable critic behavior—initially rising but quickly collapsing—which weakens the learning signal and degrades policy performance. To analyze scalability, we evaluate medium-sized ($M=20$) and large-scale ($M=50$) HetNets in Fig. 6 and Fig. 7, respectively. As the number of users increases, the HetNet becomes more heterogeneous and the state-action space grows substantially, leading to a more complex learning environment. The results show that higher HRea and VRea settings, e.g., $(12,3)$, $(16,5)$, yield better final rewards in such settings. This validates our hypothesis that larger reasoning capacity, enabled by deeper horizontal and vertical reasoning, is needed to model the increased interaction complexity.
@@
V-B2 Generalization across DRL Algorithms and Benchmark Environments
@@
Table II presents the final episodic return of five representative DRL algorithms on the HumanoidStandup-v4 task, comparing variants with and without ReaCritic. Across all methods, integrating ReaCritic leads to consistent performance improvements, with gains ranging from modest (e.g., SAC and PPO: +3.8% and +3.7%) to substantial (e.g., DDPG: +170.8%). These results demonstrate ReaCritic’s strong generalizability across different DRL backbones and task complexities. In particular, actor-critic algorithms like DDPG and TD3 benefit significantly from ReaCritic’s reasoning mechanism, which stabilizes Q-value estimation and mitigates value degradation—a common challenge in environments with long-horizon planning and high-dimensional dynamics. Even methods like A3C and PPO, which are less reliant on precise value estimation, achieve better final returns when enhanced with ReaCritic. This highlights the critic’s role in improving temporal abstraction and training stability.
@@
V-B3 Ablation Study on Horizontal Noise Exploration in HRea
@@
VI Conclusion
We proposed ReaCritic, a large reasoning transformer-based critic-model scaling framework that introduces reasoning capabilities inspired by LLMs into deep DRL for intelligent HetNet management. By combining HRea over diverse state-action sequences with VRea through stacked transformer blocks, ReaCritic overcomes the limitations of conventional shallow critics in complex, dynamic environments. Its modular design supports integration with various critic-based DRL algorithms, enabling flexible deployment across tasks. Extensive experiments demonstrate that ReaCritic enhances generalization, robustness, and convergence, offering a promising direction for intelligent and scalable network management.

### v2

@@ 5 Numerical Results
The evaluation framework is meticulously curated to encapsulate the multifaceted technical constraints of next-generation communication systems while maintaining algorithmic transparency. In particular, the custom HetNet environment replicates a dense deployment of macro and small-cell base stations, characterized by high-dimensional coupling and complex interference regimes. To encapsulate the physical layer intricacies of urban 5G/6G deployments, we incorporate Rician fading to model multipath propagation and Gaussian mobility models to account for the stochastic nature of user trajectories. These elements collectively simulate the environmental non-stationarity and unpredictability that challenge conventional learning paradigms, thereby providing a rigorous platform to evaluate the reasoning-time scaling capabilities of ReaCritic under realistic network dynamics.
5.1 Experimental Setup
5.1.1 Wireless Benchmark Environment
We construct a high-dimensional wireless environment that simulates a HetNet with $M\in\left[10,50\right]$ users. As the number of users increases, the decision space expands significantly, making per-user resource allocation highly complex and inefficient. In this multi-user scenario, we consider a network central node that makes global decisions for all users. In this setting, each user is randomly located in a normalized two-dimensional area and exhibits independent mobility modeled by Gaussian perturbations at each time step. Channel conditions follow a Rician fading model combined with log-distance path loss, with transmission distances sampled from $\left[5,10\right]$ meters. Channel gains are updated in each episode to capture temporal variation.
Each user is characterized by heterogeneous attributes, including task demand, latency requirement, computation capability, channel quality, physical location, and a one-hot encoded type indicator. At each decision step, the agent selects a five-dimensional action vector per user that determines uplink power, downlink power, bandwidth, and computation allocation. Actions are normalized and clipped to satisfy global resource constraints. The environment then computes uplink and downlink rates, total task latency, and a multi-objective reward that integrates throughput, energy efficiency, resource utilization, and latency satisfaction.
5.1.2 DRL Benchmark Environments
Furthermore, we utilize standardized benchmarks from the OpenAI Gym library to assess the generalizability of the proposed framework and to ensure experimental reproducibility. High-dimensional continuous control tasks such as HumanoidStandup-v4 are selected due to their requirement for intricate coordination across highly dependent state variables, which serves as a functional proxy for the resource allocation complexities inherent in wireless networks. By demonstrating consistent performance gains across both domain-specific wireless scenarios and established control benchmarks, we substantiate the robust architectural versatility of our dual-axis reasoning mechanism. This multifaceted evaluation confirms that the efficacy of horizontal and vertical reasoning is both practically relevant to wireless systems and theoretically generalizable to broader classes of high-dimensional optimization problems across varying levels of environmental complexity. To evaluate the general applicability of ReaCritic, we conduct experiments across two categories of benchmark environments. First, we compare against standard DRL algorithms, i.e., SAC, DDPG, A3C, TD3, and PPO, within our custom wireless resource allocation environment. These serve as strong baselines for measuring learning stability, sample efficiency, and final policy performance. Second, to assess task generalizability beyond wireless systems, we integrate ReaCritic into these DRL algorithms and test them on three representative OpenAI Gym [4] control tasks with varying complexity:
@@ 5.1.2 DRL Benchmark Environments
These environments provide diverse and challenging domains to validate ReaCritic’s ability to enhance value estimation and policy learning across different settings.
5.2 Experiments Performance Analysis
5.2.1 Performance Evaluation in HetNet with Varying User Density
@@
Figure 5: Comparison of training performance with block number to be 44, and user number to be 10 across ReaCritic-based SAC method with varying numbers of HRea steps and VRea steps, and standard SAC. The final reward ReaCritic-based SAC with different settings of HRea step and VRea step for 10 users is also included.
@@
Figure 6: Comparison of training performance with block number to be 44, and user number to be 20 across ReaCritic-based SAC method, and standard SAC. The final reward ReaCritic-based SAC with different settings of HRea step and VRea step for 20 users is also included.
@@
Figure 7: Comparison of training performance with block number to be 44, and user number to be 50 across ReaCritic-based SAC method, and standard SAC. The final reward ReaCritic-based SAC with different settings of HRea step and VRea step for 50 users is also included.
Fig. 5 compares the training performance of ReaCritic-based SAC under different HRea and VRea configurations with a small HetNet of $10$ users. Notably, the configuration $(H=1,V=1)$ illustrated in Fig. 5(a) serves as the foundational Transformer-based baseline. This setting represents a vanilla Transformer-based critic that utilizes self-attention to capture user dependencies but lacks the iterative reasoning-time scaling proposed in this work. Compared to the standard SAC baseline, ReaCritic consistently achieves higher and more stable cumulative rewards. The relationship between HRea and VRea is complementary: while HRea steps expand the reasoning breadth by generating multiple parallel tokens to explore state-action correlations, VRea steps enhance the reasoning depth through hierarchical abstraction. The results show that although the $(1,1)$ baseline outperforms standard MLP-based SAC, scaling both dimensions to configurations such as $(4,3)$ or $(8,1)$ remains necessary to fully capture the intricate coupling in wireless networks. When both HRea and VRea steps are small, e.g., $(4,3)$ or $(8,1)$, the model already outperforms standard SAC. Moreover, the critic output remains stable and well-aligned with the reward trajectory, indicating a coherent value function approximation. In contrast, standard SAC exhibits unstable critic behavior, with initial improvement followed by rapid collapse. This weakens the learning signal and degrades policy performance. To analyze scalability, we evaluate medium-sized ($M=20$) and large-scale ($M=50$) HetNets in Fig. 6 and Fig. 7, respectively. As the number of users increases, the HetNet becomes more heterogeneous and the state-action space grows substantially, leading to a more complex learning environment. The results show that higher HRea and VRea settings, e.g., $(12,3)$, $(16,5)$, yield better final rewards in such settings. The results confirm that while the basic $(1,1)$ Transformer baseline provides a stable starting point, higher configurations like $(12,3)$ or $(16,5)$ are vital for handling the increased interaction complexity in large-scale scenarios. This validates our hypothesis that larger reasoning capacity, enabled by deeper horizontal and vertical reasoning, is needed to model the increased interaction complexity.
@@
5.2.2 Generalization across DRL Algorithms and Benchmark Environments
@@
Table II presents the final episodic return of five representative DRL algorithms on the HumanoidStandup-v4 task, comparing variants with and without ReaCritic. Across all methods, integrating ReaCritic leads to consistent performance improvements, with gains ranging from modest (e.g., SAC and PPO: +3.8% and +3.7%) to substantial (e.g., DDPG: +170.8%). These results demonstrate ReaCritic’s strong generalizability across different DRL backbones and task complexities. These results demonstrate the robust generalizability of ReaCritic across different DRL backbones, validating its efficacy as a versatile enhancement framework for complex, high-dimensional tasks. In particular, actor–critic algorithms such as DDPG and TD3 gain substantial improvements from ReaCritic’s reasoning mechanism. The dual-axis reasoning architecture stabilizes Q-value estimation and effectively mitigates value degradation, a critical bottleneck in the long-horizon and highly coupled environments characteristic of large-scale wireless networks. Moreover, methods such as A3C and PPO, though less dependent on accurate value estimates, also achieve higher returns when combined with ReaCritic. This suggests that the increased reasoning capacity provided by ReaCritic enhances the critic’s ability to provide high-quality baseline signals, thereby improving training stability and exploration efficiency even in policy-gradient-based algorithms. The consistent gains across these diverse strong baselines reinforce the credibility of our approach for deployment in heterogeneous and dynamic wireless communication scenarios.
@@
5.2.3 Ablation Study on Horizontal Noise Exploration in HRea
@@
6 Conclusion
We proposed ReaCritic, a reasoning transformer-based critic-model scaling framework that introduces reasoning-like capabilities inspired by LLMs into DRL for intelligent HetNet management. By combining HRea over diverse state-action sequences with VRea through stacked transformer blocks, ReaCritic overcomes the limitations of conventional shallow critics in complex, dynamic environments. Its modular design supports integration with various critic-based DRL algorithms, enabling flexible deployment across tasks. Extensive experiments demonstrate that ReaCritic enhances generalization, robustness, and convergence, offering a promising direction for intelligent and scalable network management.